# Mean-field approach to finite-size fluctuations in the Kuramoto-Sakaguchi model

### v1

@@ Abstract
We develop an ab initio approach to describe the statistical behavior of finite-size fluctuations in the Kuramoto-Sakaguchi model. We obtain explicit expressions for the covariance function of fluctuations of the complex order parameter and determine the variance of its magnitude entirely in terms of the equation parameters. Our results rely on an explicit complex-valued formula for solutions of the Adler equation. We present analytical results for both the sub- and the super-critical case. Moreover, our framework does not require any prior knowledge about the structure of the partially synchronized state. We corroborate our results with numerical simulations of the full Kuramoto-Sakaguchi model. The proposed methodology is sufficiently general such that it can be applied to other interacting particle systems.
@@
Introduction. Synchronization is a generic phenomenon inherent in interacting oscillators that determines their ability to exhibit a wide range of complex dynamical behavior [Kuramoto84, Str2000, PikovskyRosenblumKurths01, AceBVRS2005, ArenasEtAl08, DorflerBullo14, RodriguesEtAl16]. It has been observed in a plethora of natural and engineered systems, including pace-maker cells of circadian rhythms [YamaguchiETAl03], networks of neurons [BhowmikShanahan12], chemical oscillators [KissEtAl02, TaylorEtAl09] and power grids [FilatrellaEtAl08]. The main features of synchronization are described by the paradigmatic Kuramoto-Sakaguchi (KS) model [SakK1986]
@@
for the collective behavior of a system of $N$ all-to-all coupled phase oscillators. In this model, $\omega_{i}$ are natural frequencies drawn randomly and independently from a given distribution $g(\omega)$ and $K$ and $\lambda$ are real parameters quantifying the coupling strength and the phase lag, respectively. It is often convenient to rewrite (1) in the corotating frame $\theta_{i}=\psi_{i}-\Omega t$ with a global rotation frequency $\Omega$ as
$\dot{\theta}_{i}=\omega_{i}-\Omega+K\>\mathrm{Im}\left(Z(t)e^{-i\theta_{i}}e^{-i\lambda}\right),$ (2)
which reveals that each oscillator interacts with the other oscillators only via the mean field. Importantly, the degree of synchronization is quantified by the mean field, also coined the complex order parameter,
$Z(t)=\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}e^{i\theta_{j}(t)}.$ (3)
A fully synchronized state with $\theta_{1}=\theta_{2}=\dots=\theta_{N}$ is characterized by $r(t)=|Z(t)|=1$, whereas a completely disordered (incoherent) state is characterized by $r(t)\sim\mathcal{O}(1/\sqrt{N})$. It is well known that for sufficiently large values of the coupling strength $K$ the system becomes fully synchronized, whereas for sufficiently small values of $K$ the oscillators behave independently and form an incoherent state. For intermediate values of the coupling strength, partially synchronized states are observed in which $r(t)$ fluctuates around a constant value between 0 and $1$. On a microscopic level the oscillator population splits into two groups: coherent oscillators $\mathcal{C}$, which are phase locked, and rogue oscillators $\mathcal{R}$, which drift with respect to the mean field.
In the thermodynamic limit of infinitely many oscillators each oscillator interacts with the other oscillators via a constant collective mean. This limit is well understood and described by mean-field theory [SakK1986, OmeW2012, OmeW2013, OttA2008]. In particular, the rotation frequency of the coherent oscillators $\Omega_{\infty}$ and the constant order parameter $r_{\infty}$ are determined by a self-consistency relation
@@
However, recently finite-size effects have gained much attention. Numerical experiments have lead to a better understanding of finite-size effects and provide clear evidence for their significance [Dai1994, Dai1996, CraD1999, NisTHA2012, HonCTP2015, SnyderEtAl21, YueG2024, SumJ2024]. In particular, several finite-size phenomena were observed such as a stochastic drift of oscillators which disappears in the thermodynamic limit [GiacominPoquet15, Lucon15, Got2015, Gottwald17] and emergent random chimera switching in a deterministic two-population KS model [GI25]. Capturing finite-size effects analytically is a notoriously hard problem. Relaxation rates and multi-oscillator were obtained using kinetic theory [HildebrandEtAl07, BuiceChow07, YoonEtAl15] and finite-size scaling near criticality has been successfully quantified [HongEtAl07, HonCTP2015, HonOS2016, ParkPark24]. A general scheme for analyzing fluctuations in systems of coupled phase oscillators was proposed in [KurN1987, Daido89, Dai1990]. In particular, accurate approximations were obtained in the subcritical case of the KS model (1) with $\lambda=0$ and a Lorentzian frequency distribution [Dai1990]. However, this approach relies on several approximations which do not carry over to the case of more general frequency distributions such as Gaussians as we will show (see Appendix). In this Letter we will capture finite-size effects in a first principle way and provide explicit expressions for the covariance function and the variance of the order parameter. Our focus will be the hitherto unexplored parameter range away from criticality.
Recently, it was numerically recognized that fluctuations
$\zeta(t)=\sqrt{N}(Z(t)-\langle Z(t)\rangle)$ (6)
around the thermodynamic mean $\langle Z(t)\rangle$, where $\langle\cdot\rangle$ denotes time averaging, approximately obey a Gaussian process [YueG2024, GI25]. Fluctuations in $Z(t)$, with a variance that decays as $1/N$ and vanishes in the thermodynamic limit, lead to fluctuations in other macroscopic variables, including the order parameter with fluctuations
$\delta(t)=\sqrt{N}(|Z(t)|-\langle|Z(t)|\rangle)=\sqrt{N}(r(t)-\langle r(t)\rangle).$ (7)
@@
$R(\tau)=\langle\zeta(t)\overline{\zeta(t+\tau)}\rangle,$ (8)
where the bar denotes the complex-conjugate. Once determined, this suggests to approximate the deterministic dynamics of the synchronized oscillators $\theta_{i}$, which are deterministically driven by $Z(t)$, by an effective stochastic equation for $\theta_{i}$, which is driven by a Gaussian process [Givonetal04]. Previous work [YueG2024, GI25] has shown the efficacy of such an approach, however, these works relied on approximating the general Gaussian process by an Ornstein-Uhlenbeck process the parameters of which were estimated only numerically and needed to be recalculated for each set of equation parameters. Moreover, a precise a priori knowledge of which oscillators are synchronized and which are not was required.
In this Letter, we find explicit analytical expressions for the covariance function (8) and the variance of finite-size fluctuations of the order parameter $V=\langle\delta^{2}(t)\rangle$, given entirely in terms of equation parameters $K$, $\lambda$ and $g(\omega)$. We present expressions for both the partially synchronized and the completely incoherent state. Remarkably, our expressions do not require any a priori knowledge about which oscillators partake in the synchronized cluster and which do not.
@@
Our analytical results and their efficacy in capturing finite-size fluctuations are illustrated in Figs. 1 and 2, where we compare our predictions with simulations of the full KS model (1) for a network of $1,000$ oscillators. Fig. 1 shows the covariance function $R(\tau)$ for a partially synchronized state described by (9) and for a completely incoherent state described by (11). Fig. 2(a) shows the averaged order parameter $\langle r(t)\rangle$ as a function of the coupling strength $K$. The order parameter of the full KS model (1) is very well reproduced by classical mean-field theory (4) for values of the coupling strength $K\geq K_{\rm{crit}}$. Moreover, our extension of the mean-field results to the completely incoherent state given by (12) captures the value of the order parameter for the subcritical range $K<K_{\rm{crit}}$. Fig. 2(b) shows how our analytical results for the variance of the order parameter (10), (12) reproduce the true variance well for coupling strength away from the bifurcation at $K_{\rm{crit}}$. We have checked that the results for the full KS model (1) are indistinguishable by eye when a larger network with $10,000$ oscillators is simulated. Additional numerical results for various values of the coupling strength $K$ are provided in the Appendix.
@@
Figure 2: (a) Averaged order parameter $\langle r(t)\rangle$, with theoretical predictions from classical mean-field theory (4) for $K>K_{\rm{crit}}$ and from our approach (12) for $K<K_{\rm{crit}}$, and (b) its variance $V$, with theoretical predictions (10), (12), versus coupling strength $K$ in the KS model (1) with $\lambda=\pi/4$.
Covariance function for a partially synchronized state. Our approach is based on the following prerequisites. (i) Although seeking to quantify fluctuations of $Z(t)$, we replace in Eq. (2) the mean-field $Z(t)$ by its constant thermodynamic value $Z_{\infty}$, and we move into the corotating frame
$\theta_{i}(t)=\psi_{i}(t)-\Omega_{\infty}t+\phi\quad\mbox{with some}\quad\phi\in\mathbb{R}.$ (13)
Moreover, we assume $\langle Z(t)\rangle=Z_{\infty}$. (ii) We use the well-known fact (see Appendix for details) that the probability density of the solution to Eq. (2) with $Z(t)=Z_{\infty}$ is given by the Ott-Antonson ansatz [OttA2008]
@@
Then, for every non-negative integer $n$ we have
@@
$Z_{\infty}=\int_{-\infty}^{\infty}d\omega\int_{0}^{2\pi}f_{\phi}(\theta,\omega)e^{i\theta}d\theta=-r_{\infty}ie^{i\lambda}e^{i\phi}.$
Moreover, due to the ergodicity property, for any $2\pi$-periodic function $F(\theta)$ we have
@@
(iii) Finally, we establish an explicit solution of the thermodynamic limit of (2) with constant driver $Z(t)=Z_{\infty}$, the so called Adler equation. Introducing $\Delta_{j}=\omega_{j}-\Omega_{\infty}$ and $H_{\infty}=KZ_{\infty}e^{-i\lambda}$, solutions of the Adler equation $\theta_{j}(t)$ with initial conditions $e^{i\theta_{j}(0)}$ are given for $|\Delta_{j}|\neq|H_{\infty}|$ by the explicit formula
@@
$\xi=\left\{\begin{array}[]{lcl}-i\sqrt{|H|^{2}-\Delta^{2}}&\mbox{ for }&|\Delta|\leq|H|,\\[5.69054pt] \sqrt{\Delta^{2}-|H|^{2}}&\mbox{ for }&|\Delta|>|H|.\end{array}\right$. (20)
Note that for $|\Delta_{j}|<|H_{\infty}|$, the $j$th oscillator is coherent and the dynamics of $\theta_{j}(t)$ converges to a fixed point as $t\to\infty$, while for $|\Delta_{j}|>|H_{\infty}|$, the oscillator is incoherent with underlying periodic dynamics. Details on the derivation of (18), (19) are given in the Appendix.
@@
$R_{\mathrm{ps}}(\tau)$ $=$ $N\langle Z(t)\overline{Z(t+\tau)}\rangle-N|Z_{\infty}|^{2}$ (21)
@@
It is easy to see that if $j\in\mathcal{C}$ or $k\in\mathcal{C}$, the summands can be written as
@@
Moreover, assuming that the variables $e^{i\theta_{j}(t)}$ and $e^{i\theta_{k}(t)}$ with $j,k\in\mathcal{R}$ and $j\neq k$ are independent, and therefore identity (22) holds, the sum in (21) can be written as
@@
Substitution in (21) yields, upon using (6),
$R_{\mathrm{ps}}(\tau)=\sum\limits_{j\in\mathcal{R}}\langle e^{i\theta_{j}(t)}e^{-i\theta_{j}(t+\tau)}\rangle-\sum\limits_{j\in\mathcal{R}}\left|\langle e^{i\theta_{j}(t)}\rangle\right|^{2}.$ (23)
@@
(Note that for the second correlation, we used identity (16) after expanding the solution (18)–(19) in a geometric series. Details of the calculations can be found in the Appendix.) Altogether, this yields
@@
After some algebraic manipulations (see Appendix), the covariance function (24) simplifies to our expression (9). Note that if $|\omega-\Omega|\leq|H_{\infty}|=Kr_{\infty}$, $|z_{\phi}|=1$. Therefore such values $\omega$ do not contribute to the integral $R_{\mathrm{ps}}(\tau)$. This has the desirable consequence that in order to evaluate the covariance function, no prior knowledge is required about which oscillators are and which are not synchronized.
@@
(ii) The global order parameter shows no collective oscillations, and $Z(t)=0$ as $N\to\infty$. This implies that the dynamics of each oscillator is approximately described by $\dot{\psi}_{i}=\omega_{i}$, and therefore $\psi_{i}(t)=\omega_{i}t+\phi$ with $\phi=\psi_{i}(0)$.
The covariance function for fluctuations $\zeta_{\mathrm{incoh}}(t)$ of this incoherent state is then given by
@@
For randomly chosen initial conditions $\psi_{i}(0)$ the variables $e^{i\psi_{j}(t)}$ and $e^{i\psi_{k}(t)}$ with $j\neq k$ are independent, which implies again an identity of the form (22), and we obtain
@@
$\langle e^{i\psi_{j}(t)}\rangle=\int_{0}^{2\pi}e^{i\psi}\frac{\displaystyle f_{\mathrm{incoh}}(\psi,\omega,t)}{\displaystyle g(\omega)}d\psi=0$ (26)
@@
Finally, in the large-$N$ limit we can replace averaging over indices $j$ by averaging over the distribution $g(\omega)$ and arrive at our expression for the covariance function $R_{\rm{incoh}}$ (11). Note that due to the normalization condition for $g(\omega)$, we have $\langle\left|\zeta_{\mathrm{incoh}}(t)\right|^{2}\rangle=R_{\mathrm{incoh}}(0)=1$.
@@
$Z(t)=\langle Z(t)\rangle+\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\zeta(t)=Z_{\infty}+\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\zeta(t)$
@@
$|Z(t)|^{2}=|Z_{\infty}|^{2}+\frac{\displaystyle 2}{\displaystyle\sqrt{N}}\mathrm{Re}\>\left(\overline{Z_{\infty}}\zeta(t)\right)+\frac{\displaystyle 1}{\displaystyle N}|\zeta(t)|^{2},$ (27)
@@
$|Z(t)|=|Z_{\infty}|+\frac{\displaystyle 1}{\displaystyle\sqrt{N}|Z_{\infty}|}\mathrm{Re}\>\left(\overline{Z_{\infty}}\zeta(t)\right)+\mathcal{O}\left(\frac{\displaystyle 1}{\displaystyle N}\right).$
@@
where $\tilde{R}_{\rm{ps}}(t)$ denotes the pseudo-covariance $\tilde{R}_{\rm{ps}}(t)=\langle\zeta(t)\zeta(t+\tau)\rangle$. Employing calculations similar to those for $R_{\mathrm{ps}}(t)$ we can show that $\tilde{R}_{\mathrm{ps}}(\tau)=0$ for all $\tau\geq 0$. The vanishing of the pseudo-covariance can be intuitively understood by evoking a symmetry between the real and imaginary parts of the fluctuations $\zeta$ and their independence. Using $\tilde{R}_{\mathrm{ps}}(\tau)=0$ for all $\tau\geq 0$, leads to our result (10).
Variance of the order parameter for a completely incoherent state. Consider fluctuations $\zeta_{\mathrm{incoh}}(t)$ of the complex order parameter in the incoherent state. In the large-$N$ limit, for a completely incoherent state we have $\langle Z(t)\rangle=0$, and (27) reduces to
$|Z(t)|=\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\left|\zeta_{\mathrm{incoh}}(t)\right|.$
Using the definition of fluctuations (6), we write
@@
where $\langle\cdot\rangle_{\mathrm{ens}}$ denotes the ensemble average. Similarly, $\sin\psi_{j}(t)\sim\mathcal{N}(0,1/2)$. Moreover, it is easy to verify that the random variables $\cos\psi_{j}(t)$ and $\sin\psi_{j}(t)$ are uncorrelated, since
@@
Hence $\mathrm{Re}\>\zeta_{\mathrm{incoh}}(t)\sim\mathcal{N}(0,1/2)$ and $\mathrm{Im}\>\>\zeta_{\mathrm{incoh}}(t)\sim\mathcal{N}(0,1/2)$, and $\left|\zeta_{\mathrm{incoh}}(t)\right|$ has a Rayleigh distribution with expected value
@@
Discussion and outlook. In this Letter, we have proposed a fully analytical approach to capture the statistical behavior of finite-size fluctuations in the KS model (1). In particular, our approach provides

### v2

@@ Abstract
We develop an ab initio approach to describe the statistical behavior of finite-size fluctuations in the deterministic Kuramoto-Sakaguchi model. We obtain explicit expressions for the covariance function of fluctuations of the complex order parameter and determine the variance of its magnitude entirely in terms of the equation parameters. Our results rely on an explicit complex-valued formula for the solution of the Adler equation. We present analytical results for both the sub- and the super-critical case. Moreover, our framework does not require any prior knowledge about the structure of the partially synchronized state. We corroborate our results with numerical simulations of the full Kuramoto-Sakaguchi model. The proposed methodology is sufficiently general such that it can be applied to other interacting particle systems.
@@
Introduction. Synchronization is a generic phenomenon inherent in interacting oscillators that determines their ability to exhibit a wide range of complex dynamical behavior [27, 48, 41, 1, 2, 13, 44]. It has been observed in a plethora of natural and engineered systems, including pace-maker cells of circadian rhythms [53], networks of neurons [3], chemical oscillators [25, 50] and power grids [15]. The main features of synchronization are captured by the paradigmatic Kuramoto-Sakaguchi (KS) model [45]
@@
for the collective behavior of a system of $N$ all-to-all coupled phase oscillators. In this model, $\omega_{i}$ are natural frequencies drawn randomly and independently from a given distribution $g(\omega)$ and $K$ and $\lambda$ are real parameters quantifying the coupling strength and the phase lag, respectively. It is often convenient to rewrite Eq. (1) as
$\dot{\psi_{i}}=\omega_{i}+K\>\mathrm{Im}\left(Z(t)e^{-i\psi_{i}}e^{-i\lambda}\right),$
which reveals that each oscillator interacts with the other oscillators only via the mean field
$Z(t)=\frac{\displaystyle 1}{\displaystyle N}\sum\limits_{j=1}^{N}e^{i\psi_{j}(t)},$ (2)
which is also coined the complex order parameter and quantifies the degree of synchronization.
A fully synchronized state with $\theta_{1}=\theta_{2}=\dots=\theta_{N}$ is characterized by $r(t)=|Z(t)|=1$, whereas a completely disordered (incoherent) state is characterized by $r(t)=\mathcal{O}(1/\sqrt{N})$. It is well known that for sufficiently large values of the coupling strength $K$ the system becomes fully synchronized, whereas for sufficiently small values of $K$ the oscillators behave independently and form an incoherent state. For intermediate values of the coupling strength, partially synchronized states are observed in which $r(t)$ fluctuates around a constant value between 0 and $1$. On a microscopic level the oscillator population splits into two groups: coherent oscillators $\mathcal{C}$, which are phase locked, and rogue oscillators $\mathcal{R}$, which drift with respect to the mean field.
In the thermodynamic limit of infinitely many oscillators each oscillator interacts with the other oscillators via a constant collective mean. This limit is well understood and described by mean-field theory [45, 33, 34, 35]. In particular, the rotation frequency of the coherent oscillators $\Omega_{\infty}$ and the constant order parameter $r_{\infty}$ are determined by a self-consistency relation, first derived in [45] and later formulated in a more convenient complex form in [33],
@@
However, recently finite-size effects have gained much attention. Numerical experiments have led to a better understanding of finite-size effects and provided clear evidence for their significance [9, 10, 7, 32, 22, 47, 55, 49]. In particular, several finite-size phenomena were observed such as a stochastic drift of oscillators which disappears in the thermodynamic limit [16, 30, 18, 19] and emergent random chimera switching in a deterministic two-population KS model [24]. Capturing finite-size effects analytically is a notoriously hard problem. Relaxation rates and pair correlation functions were obtained using kinetic theory [20, 5, 54] and the finite-size scaling near criticality has been successfully quantified [23, 22, 21, 37]. Recently, substantial progress has been made for stochastic systems of coupled oscillators [52, 4]. However, analytical methods for deterministic systems still remain underdeveloped. A general scheme for analyzing fluctuations in deterministic systems of coupled phase oscillators was proposed in [26, 11, 8]. In particular, accurate approximations were obtained in the subcritical case of the KS model (1) with $\lambda=0$ and a Lorentzian frequency distribution [8]. However, this approach relies on several approximations which do not carry over to the case of more general frequency distributions such as Gaussians as we will show (see Appendix K). In this Letter we will capture finite-size effects in a first principle way and provide explicit expressions for the covariance function and the variance of the order parameter.
Our theory is based on the assumption of an underlying central limit theorem describing fluctuations around the mean-field theory limit. This implies that fluctuations in the deterministic KS model decay with $1/\sqrt{N}$ and constitute a Gaussian process determined entirely by its mean and covariance function (see Appendix F for numerical evidence for this assumption). This points to the range of validity of our theory: we require sufficiently large system size $N$ in order to allow for the validity of the central limit theorem.
Specifically, we consider the modified order parameter
$Z_{\mathrm{mod}}(t)=Z(t)\frac{\overline{Z_{\mathrm{c}}(t)}}{|Z_{\mathrm{c}}(t)|}\quad\mbox{with}\quad Z_{\mathrm{c}}(t)=\frac{1}{N}\sum\limits_{j\in\mathcal{C}}e^{i\psi_{j}(t)},$
where the bar denotes the complex-conjugate. Compared to the original order parameter (2), the modified order parameter is represented in the frame of reference moving with the synchronized cluster. This ensures that, contrary to $Z(t)$, the modified order parameter does not exhibit phase diffusion [42, 40] (see Appendix I). Recently, it was numerically recognized that fluctuations
$\zeta(t)=\sqrt{N}(Z_{\mathrm{mod}}(t)-\langle Z_{\mathrm{mod}}(t)\rangle)$ (5)
around the thermodynamic mean $\langle Z_{\mathrm{mod}}(t)\rangle$, where $\langle\cdot\rangle$ denotes time averaging, can be approximated by a Gaussian process [55, 24]. Fluctuations in $Z_{\mathrm{mod}}(t)$, with a variance that decays as $1/N$ and vanishes in the thermodynamic limit, lead to fluctuations in other macroscopic variables, including the order parameter with fluctuations
$\delta(t)=\sqrt{N}(|Z_{\mathrm{mod}}(t)|-\langle|Z_{\mathrm{mod}}(t)|\rangle)=\sqrt{N}(r(t)-\langle r(t)\rangle).$ (6)
@@
$R(\tau)=\langle\zeta(t)\overline{\zeta(t+\tau)}\rangle.$ (7)
Previous work [55, 24] has shown the efficacy of such an approach, however, these works relied on approximating the general Gaussian process by an Ornstein-Uhlenbeck process the parameters of which were estimated only numerically and needed to be recalculated for each set of equation parameters. Moreover, a precise a priori knowledge of which oscillators are synchronized and which are not was required.
In this Letter, we find explicit analytical expressions for the covariance function (7) and the variance of finite-size fluctuations of the order parameter $V=\langle\delta^{2}(t)\rangle$, given entirely in terms of equation parameters $K$, $\lambda$ and $g(\omega)$. We present expressions for both the partially synchronized and the completely incoherent state. Remarkably, our expressions do not require any a priori knowledge about which oscillators partake in the synchronized cluster and which do not. This opens up the way to perform stochastic model reduction for deterministic systems of coupled oscillators where the drift and diffusion coefficients are given entirely in terms of the microscopic equation parameters. Such reductions have so far been restricted to multi-scale systems [17].
@@
Our analytical results and their efficacy in capturing finite-size fluctuations are illustrated in Figs. 1 and 2, where we compare our predictions with simulations of the full KS model (1) for a network of $1,000$ oscillators with a Gaussian natural frequency distribution $g(\omega)$ with unit-variance. Fig. 1 shows the covariance function $R(\tau)$ for a partially synchronized state described by (8) and for a completely incoherent state described by (10). Fig. 2(a) shows the averaged order parameter $\langle r(t)\rangle$ as a function of the coupling strength $K$. The order parameter of the full KS model (1) is very well reproduced by classical mean-field theory (3) for values of the coupling strength $K\geq K_{\rm{crit}}$. Moreover, our extension of the mean-field results to the completely incoherent state given by (11) reproduces the order parameter for the subcritical range $K<K_{\rm{crit}}$. Fig. 2(b) shows that the variance of the order parameter is well captured by our analytical expressions (9) and (11) for coupling strength away from the bifurcation at $K_{\rm{crit}}$. Additional numerical results and the description of our numerical setup are provided in Appendix H and Appendix E, respectively.
@@
Figure 2: (a) Averaged order parameter $\langle r(t)\rangle$, with theoretical predictions from classical mean-field theory (3) for $K>K_{\rm{crit}}\approx 1.92$ and from our approach (11) for $K<K_{\rm{crit}}$, and (b) its variance $V$, with theoretical predictions (9), (11), versus coupling strength $K$ in the KS model (1) with $\lambda=\pi/4$.
Covariance function for a partially synchronized state. Our approach is based on the following prerequisites. (i) Moving into the corotating frame $\theta_{i}(t)=\psi_{i}(t)-\arg Z_{\mathrm{c}}(t)$ and neglecting the assumed Gaussian fluctuations for $N\to\infty$, Eq. (1) becomes the so called Adler equation
$\dot{\theta}_{i}=\omega_{i}-\Omega_{\infty}+K\>\mathrm{Im}\left(Z_{\infty}e^{-i\theta_{i}}e^{-i\lambda}\right).$ (12)
(ii) We use the well-known fact (see Appendix A for details) that the probability density of the solution of the Adler equation (12) is given by the Ott-Antonson ansatz [35]
@@
It follows directly from (13) that for every non-negative integer $n$ we have
@@
$Z_{\infty}=\int_{-\infty}^{\infty}d\omega\int_{0}^{2\pi}f_{\phi}(\theta,\omega)e^{i\theta}d\theta=-r_{\infty}ie^{i\lambda}e^{i\phi}.$ (16)
Moreover, if $\theta_{i}(t)$ is a solution of the scalar Adler equation (12), then for any $2\pi$-periodic function $F(\theta)$ we have the ergodicity property
@@
(iii) Finally, we establish an explicit solution of the Adler equation (12). Introducing $\Delta_{j}=\omega_{j}-\Omega_{\infty}$ and $H_{\infty}=KZ_{\infty}e^{-i\lambda}$, solutions of the Adler equation $\theta_{j}(t)$ with initial conditions $\theta_{j}(0)$ are given for $|\Delta_{j}|\neq|H_{\infty}|$ by
@@
$\xi=\left\{\begin{array}[]{lcl}-i\sqrt{|H|^{2}-\Delta^{2}}&\mbox{ for }&|\Delta|\leq|H|,\\[5.69054pt] \hphantom{-i}\sqrt{\Delta^{2}-|H|^{2}}&\mbox{ for }&|\Delta|>|H|.\end{array}\right$. (20)
Note that for $|\Delta_{j}|<|H_{\infty}|$, the $j$th oscillator is coherent and the dynamics of $\theta_{j}(t)$ converges to a fixed point as $t\to\infty$, while for $|\Delta_{j}|>|H_{\infty}|$, the oscillator is incoherent with underlying periodic dynamics. Details on the derivation of (18), (19) are provided in Appendix B.
@@
$R_{\mathrm{ps}}(\tau)$ $=$ $N\langle Z_{\mathrm{mod}}(t)\overline{Z_{\mathrm{mod}}(t+\tau)}\rangle-N|Z_{\infty}|^{2}$ (21)
@@
If either $j\in\mathcal{C}$ or $k\in\mathcal{C}$ the phase asymptotically approaches a fixed point and the summands can be written as
@@
Moreover, the independence identity (22) also holds if $j,k\in\mathcal{R}$ and $\sqrt{\Delta_{j}^{2}-|H_{\infty}|^{2}}$ and $\sqrt{\Delta_{k}^{2}-|H_{\infty}|^{2}}$ are incommensurable (see Appendix J), and hence generically the sum in (21) can be written as
@@
Substitution into (21) yields, upon using $\langle Z_{\mathrm{mod}}(t)\rangle=Z_{\infty}$,
$R_{\mathrm{ps}}(\tau)=\frac{1}{N}\sum\limits_{j\in\mathcal{R}}\langle e^{i\theta_{j}(t)}e^{-i\theta_{j}(t+\tau)}\rangle-\frac{1}{N}\sum\limits_{j\in\mathcal{R}}\left|\langle e^{i\theta_{j}(t)}\rangle\right|^{2}.$ (23)
@@
(Note that for the second correlation, we used identity (15) after expanding the solution (18)–(19) in a geometric series. Details of the calculations can be found in Appendix B.) Altogether, this yields
@@
After some algebraic manipulations (see Appendix C), the covariance function (24) simplifies to our expression (8). Note that if $|\omega-\Omega_{\infty}|\leq|H_{\infty}|=Kr_{\infty}$, $|z_{\phi}|=1$. Therefore such a set of $\omega$ does not contribute to the integral $R_{\mathrm{ps}}(\tau)$. This has the desirable consequence that in order to evaluate the covariance function, no prior knowledge is required about which oscillators are and which are not synchronized.
@@
(ii) The global order parameter shows no collective oscillations, and $Z(t)=0$ as $N\to\infty$. This implies that the dynamics of each oscillator is approximately described by $\dot{\psi}_{i}=\omega_{i}$, and therefore $\psi_{i}(t)=\omega_{i}t+\psi_{i}(0)$.
Consider the fluctuations
$\zeta_{\mathrm{incoh}}(t)=\sqrt{N}(Z(t)-\langle Z(t)\rangle)$
of the complex order parameter in this incoherent state. The corresponding covariance function is then given by
@@
For different incommensurable frequencies $\omega_{j}$ and $\omega_{k}$, the variables $e^{i\psi_{j}(t)}$ and $e^{i\psi_{k}(t)}$ are independent, which implies again an identity of the form (22), and we obtain
@@
$\langle e^{i\psi_{j}(t)}\rangle=\int_{0}^{2\pi}e^{i\psi}\frac{\displaystyle f_{\mathrm{incoh}}(\psi,\omega_{j},t)}{\displaystyle g(\omega_{j})}d\psi=0$ (26)
@@
Finally, in the large-$N$ limit we can replace averaging over indices $j$ by averaging over the distribution $g(\omega)$ and arrive at our expression (10) for the covariance function $R_{\rm{incoh}}(\tau)$. Note that due to the normalization condition for $g(\omega)$, we have $\langle\left|\zeta_{\mathrm{incoh}}(t)\right|^{2}\rangle=R_{\mathrm{incoh}}(0)=1$.
@@
$Z_{\mathrm{mod}}(t)=\langle Z_{\mathrm{mod}}(t)\rangle+\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\zeta(t)=Z_{\infty}+\frac{\displaystyle 1}{\displaystyle\sqrt{N}}\zeta(t)$
@@
$|Z_{\mathrm{mod}}(t)|^{2}=|Z_{\infty}|^{2}+\frac{\displaystyle 2}{\displaystyle\sqrt{N}}\mathrm{Re}\>\left(\overline{Z_{\infty}}\zeta(t)\right)+\frac{\displaystyle 1}{\displaystyle N}|\zeta(t)|^{2},$
@@
$|Z_{\mathrm{mod}}(t)|=|Z_{\infty}|+\frac{\displaystyle 1}{\displaystyle\sqrt{N}|Z_{\infty}|}\mathrm{Re}\>\left(\overline{Z_{\infty}}\zeta(t)\right)+\mathcal{O}\left(\frac{\displaystyle 1}{\displaystyle N}\right).$
@@
where $\tilde{R}_{\rm{ps}}(t)$ denotes the pseudo-covariance $\tilde{R}_{\rm{ps}}(t)=\langle\zeta(t)\zeta(t+\tau)\rangle$. Employing calculations similar to those for $R_{\mathrm{ps}}(t)$ (see Appendix D for details), we can show that $\tilde{R}_{\mathrm{ps}}(\tau)=0$ for all $\tau\geq 0$. The vanishing of the pseudo-covariance can be intuitively understood by evoking a symmetry between the real and imaginary parts of the fluctuations $\zeta$ and their independence. Using $\tilde{R}_{\mathrm{ps}}(\tau)=0$ for all $\tau\geq 0$, leads to our result (9).
Variance of the order parameter for a completely incoherent state. In the large-$N$ limit, for a completely incoherent state we have $\langle Z(t)\rangle=0$, and therefore $|Z(t)|=|\zeta_{\mathrm{incoh}}(t)|/\sqrt{N}.$ Using the definition of $\zeta_{\mathrm{incoh}}(t)$, we write
@@
where $\langle\cdot\rangle_{\mathrm{ens}}$ denotes the ensemble average. Similarly, $\sin\psi_{j}(t)\sim\mathcal{N}(0,1/2)$. Moreover, the random variables $\cos\psi_{j}(t)$ and $\sin\psi_{j}(t)$ are uncorrelated since
@@
Hence $\mathrm{Re}\>(\zeta_{\mathrm{incoh}}(t))\sim\mathcal{N}(0,1/2)$ and $\mathrm{Im}\>(\zeta_{\mathrm{incoh}}(t))\sim\mathcal{N}(0,1/2)$, and $\left|\zeta_{\mathrm{incoh}}(t)\right|$ has a Rayleigh distribution with expected value
@@
Discussion and outlook. In this Letter, we have proposed a fully analytical approach to capture the statistical behavior of finite-size fluctuations in the deterministic KS model (1). In particular, our approach provides